# Scaled Block Vecchia Approximation for High-Dimensional Gaussian Process Emulation on GPUs

### v2

@@ 7. Performance and Energy Evaluation
(c) Runtime on GH200 superchip (CPU+GPU).
(d) Performance on GH200 superchip (GPU only).
@@
All following experiments set the default $bs=100$, 10-dimensional input $0,1]^{10}$, and only consider the estimation stage, which is the most time-consuming part in the whole lifetime of emulations; thus, $m$ represents the number of nearest neighbors for estimations and uses Maximum Likelihood Estimation (MLE). Here, we set the iterations to 500, considering that the average number of optimization rounds is 500 in benchmark datasets.
@@ 7.1. Single-Node Performance
Figure 8 illustrates the runtime and GPU throughput performance of the SBV and SV methods on two nodes: an NVIDIA A100 with its host CPU, and the GH200 Superchip. Subfigures (a) and (c) show that SBV consistently outperforms SV in total execution time across different values of $m$, demonstrating superior scalability as the dataset size increases. As expected, larger values of $m$ increase the total runtime due to higher computational cost per iteration; however, they also lead to significantly better approximation and prediction accuracy, as shown in the accuracy section.
@@ 7.2. Scalability Across Multiple Nodes
Figure 9 demonstrates the weak and strong scaling performance of the distributed SBV implementation on NVIDIA A100 and GH200 architectures, scaling up to 512 GPUs. In weak scaling (Subfigures (a) and (c)), both systems maintain high parallel efficiency ($PE$) as the problem size increases proportionally with the number of GPUs. The drops in the larger scale are derived from the CPU computation, specifically Step 2 in the Algorithm 4, which can be reduced using a smaller $\lambda$ in Eq 7. Specifically, the candidate count does not scale proportionally and retains some stochastic variability, introducing non-uniform load balancing across tasks. A gradually decreasing $\lambda$ is recommended in the NNS as the GPU count increases.
In strong scaling (Subfigures (b) and (d)), total execution time decreases nearly linearly with increasing GPU count, and parallel efficiency exceeds ideal scaling ($PE>1.0$) in several cases due to less nearest neighbor candidates in the Algorithm 4, where $\lambda$ can be increased accordingly to ensure there is enough candidate for NNS. The drops in the larger GPU count originate from the overhead of multiple function calls, where the 500 iterations are executed and their accumulation cannot be hidden as the computational task is light for each GPU, where each iteration only costs around 30 milliseconds.
In both scaling experiments, the influence of neighborhood size $m_{\text{est}}$ is also evident: larger values, such as $m=400$, increase the computational workload per point and runtime, but result in better GPU throughput and overall hardware utilization. Smaller values, such as $m=100$, result in faster runtimes but less efficient GPU utilization, especially on the GH200. Mid-range values (e.g., $m=200$) balance performance and efficiency.
A detailed breakdown shows that GPU compute time dominates execution, accounting for over 80% of the total runtime in large problem sizes. However, the remaining CPU-side computation introduces variations in scaling. On A100, limited CPU resources (32 cores per task) contribute to a gradual drop in $PE$, whereas the 72-core CPU on GH200 helps better hide this overhead. This difference comes from the role of the CPU in maintaining a globally sorted array for block reordering and candidate filtering, as in Equation (2).
@@
Figure 9. Weak and strong scaling of SBV were evaluated on up to 512 GPUs, using 500 MLE iterations on two architectures: AMD EPYC with NVIDIA A100 (40 GB, up to 128M points) and the NVIDIA GH200 superchip (96 GB, up to 2.56B points).
@@ 8. Conclusion
This paper introduced the Scaled Block Vecchia (SBV) approximation, a novel framework for scalable GP emulation tailored for high-dimensional datasets on distributed GPU systems. By integrating anisotropic input scaling with block-based conditioning and leveraging batched GPU kernels, SBV substantially reduces memory and computational costs while maintaining high predictive accuracy. To our knowledge, this is the first distributed implementation of a Vecchia-based GP approximation. Through extensive experiments on synthetic and real-world data, including the satellite drag benchmark and MetaRVM epidemiological model, we demonstrate the scalability up to 512 GPUs and the ability to handle up to 2.52.5B points with energy efficiency and state-of-the-art accuracy.
@@ Acknowledgments
This work was supported and funded by King Abdullah University of Science and Technology (KAUST). We gratefully acknowledge Ahmad Abdelfattah from the University of Tennessee, Knoxville (UTK), for his support and assistance in optimizing the use of the MAGMA library. We also thank the KAUST Supercomputing Laboratory (KSL) team for providing access to the Ibex cluster for initial runs and experiments on available GPUs. We further acknowledge the computing time provided by the John von Neumann Institute for Computing (NIC) on the supercomputer JURECA-DC. We are grateful for access to JEDI and JUPITER, which is provided through the JUPITER Research and Early Access Program (JUREAP).

### v3

@@ 8. Performance and Energy Evaluation
(c) Runtime on GH200 superchip
(CPU+GPU).
(d) Performance on GH200 superchip
(GPU only).
@@
All following experiments set the default $bs=100$, 10-dimensional input $0,1]^{10}$, and only consider the estimation stage, which is the most time-consuming part in the whole lifetime of emulations; thus, $m$ represents the number of nearest neighbors for estimations and uses Maximum Likelihood Estimation (MLE). Here, we set the number of iterations to 500, as the average number of optimization rounds is 500 on benchmark datasets.
@@ 8.1. Single-Node Performance
Figure 8 illustrates the runtime and GPU throughput performance of the SBV and SV methods on two nodes: an NVIDIA A100 with its host CPU, and the GH200 Superchip. Subfigures (a) and (c) show that SBV consistently outperforms SV in total execution time across different values of $m$, demonstrating superior scalability as the dataset size increases. As expected, larger values of $m$ increase the total runtime due to the higher computational cost per iteration; however, they also yield significantly better approximation and prediction accuracy, as shown in the accuracy section.
@@
While the GH200 offers substantially higher theoretical FP64 peak performance compared to the A100, the SBV workload is not exclusively compute-bound. The primary kernels involve batched Cholesky factorizations and moderate-sized GEMM/TRSM operations on blocks with $bs\approx 100$ and $m\leq 400$. For these matrix dimensions, arithmetic intensity remains moderate, and performance is limited in part by memory bandwidth and batched-kernel occupancy rather than by peak FP64 throughput. As a result, the observed performance gains correspond more closely to increases in effective HBM bandwidth than to theoretical peak FLOP ratios. This pattern suggests that SBV operates within a mixed memory-and-compute regime, rather than failing to utilize the capabilities of the newer architecture.
@@ 8.2. Scalability Across Multiple Nodes
Figure 9 demonstrates the weak and strong scaling performance of the distributed SBV implementation on NVIDIA A100 and GH200 architectures, scaling up to 512 GPUs. In weak scaling (Subfigures (a) and (c)), both systems maintain high parallel efficiency ($PE$) as the problem size increases proportionally with the number of GPUs. The drops in the larger scale are derived from the CPU computation, specifically Step 2 in the Algorithm 4, which can be reduced using a smaller $\lambda$ in Eq 7. Specifically, the candidate count does not scale proportionally and retains some stochastic variability, introducing non-uniform load balancing across tasks. A gradually decreasing $\lambda$ is recommended in the NNS as the number of GPUs increases.
In strong scaling (Subfigures (b) and (d)), total execution time decreases nearly linearly with increasing GPU count, and parallel efficiency exceeds ideal scaling ($PE>1.0$) in several cases due to less nearest neighbor candidates in the Algorithm 4, where $\lambda$ can be increased accordingly to ensure there is enough candidate for NNS. The drops in the larger GPU count originate from the overhead of multiple function calls, where the 500 iterations are executed, and their accumulation cannot be hidden as the computational task is light for each GPU, where each iteration only costs around 30 milliseconds.
In both scaling experiments, the influence of the neighborhood size $m_{\text{est}}$ is also evident: larger values, such as $m=400$, increase the computational workload per point and runtime, but yield higher GPU throughput and overall hardware utilization. Smaller values, such as $m=100$, result in faster runtimes but less efficient GPU utilization, especially on the GH200. Mid-range values (e.g., $m=200$) balance performance and efficiency.
A detailed breakdown shows that GPU compute time dominates execution, accounting for over 80% of the total runtime in large problem sizes. However, the remaining CPU-side computation introduces scaling variations. On A100, limited CPU resources (32 cores per task) lead to a gradual drop in $PE$, whereas the 72-core CPU on GH200 better hides this overhead. This difference arises from the CPU’s role in maintaining a globally sorted array for block reordering and candidate filtering, as in Equation (2).
@@
Figure 9. Weak and strong scaling of SBV were evaluated on up to 512 GPUs, using 500 MLE iterations on two architectures: AMD EPYC with NVIDIA A100 (40 GB, up to 128M points) and the NVIDIA GH200 superchip (96 GB, up to 2.56B points); $T_{1}$ and $T_{p}$ represent the consumed time for single and $p$ GPUs, respectively, in the Parallel Efficiency (PE) = $\frac{T_{1}}{n\,T_{p}}$.
@@ 9. Conclusion
This paper introduced the Scaled Block Vecchia (SBV) approximation, a novel framework for scalable GP emulation tailored for high-dimensional datasets on distributed GPU systems. By integrating anisotropic input scaling with block-based conditioning and leveraging batched GPU kernels, SBV substantially reduces memory and computational costs while maintaining high predictive accuracy. To our knowledge, this is the first distributed implementation of a Vecchia-based GP approximation. Through extensive experiments on synthetic and real-world data, including the satellite drag benchmark and the MetaRVM epidemiological model, we demonstrate scalability up to $512$ GPUs and the ability to handle up to 2.52.5B points, achieving energy efficiency and state-of-the-art accuracy.
@@ Acknowledgments
This work was supported by King Abdullah University of Science and Technology (KAUST). We gratefully acknowledge Ahmad Abdelfattah from the University of Tennessee, Knoxville (UTK), for his support and assistance in optimizing the use of the MAGMA library. We also thank the KAUST Supercomputing Laboratory (KSL) team for providing access to the Ibex cluster for initial runs and GPU experiments. We further acknowledge the computing time provided by the John von Neumann Institute for Computing (NIC) on the supercomputer JURECA-DC. This project received access to JEDI, a preparation system for the JUPITER supercomputer, and JUPITER through the JUPITER Research and Early Access Program (JUREAP). JUPITER is funded by the EuroHPC Joint Undertaking, the German Federal Ministry of Research, Technology, and Space, and the Ministry of Culture and Science of the German state of North Rhine-Westphalia.